Type or pas2. Let the population regression model between a dependent variable y and an independent variable is given by y= Bo+ B1 x x+(c) Let B i Show that if B1 0, then E(30|x1,X2,...,xn) = Bo. Is it true that if E(Bola1, x2, ... ,xn) = Bo then 1 = 0? Explai
Answer #1

We can write Bo = xi = E(Bo+B1X; + u:))} = Bo + B28 +ū. Since E(u;) = 0, we have E(ū) = 0 so that E(BolX1,.., Xn) = Bo + B12

C) Since, Bo = * = E(Bo+B1x; + u;) = Bo + B17 +ū, we get Var (Polx...,xn) = Var(a) = Now Bo =ý-Bit, B1 == {(x; – 1)(): – 7) /

For any query in above, comment.